How many number of five digits can be formed without repetation if 2,3 and5 always occur in each number

Grade:10

1 Answers

Arun
6 years ago
Let's think of 2 cases. 
1. A formed number 
has 2, 3 or 5 in 10 thousands' place-
Out of remaining 4 places, 2 places are fixed. Now, out of remaining 2 places, one can be filled with 7 digits ( except 2, 3 & 5) & other with digits ( One which fills first unknown place is excluded.). 
Position of these 4 digits can be reversed in 4! ways. 
But doing this, same condition occurs twice. E.g. We have selected (6,7) & (7,6) as well.
So, we have to divide the total number by 2. 
As 5th place can be filled with 2, 3 or 5, we have
Numbers formed by condition 1 = 4!*7*6*3/2=63*4!
2. A formed number has no 2, 3 or 5 in its 5th place-
5th place can be filled by digits  
(excluding 2,3,5 & 0). If 0 is in 5th place, it will be a 4 digited number. Remaining unknown place can be filled by digits (excluding 2,3,5 & one which is in 5th place). 
If we keep 5th place fixed, other 4 digits can be reversed in 4! ways. 
Numbers formed by condition 2 = 4!*6*6
Total numbers = 4!*63+4!*36
                            =24×99=2376.